You sell one Xerox June 60 call contract and sell one Xerox June 60 put contract. The call premium is $5 and the put premium is $3.

Your strategy is called:

  1. a short straddle.

  2. a long straddle.

  3. a horizontal straddle.

  4. a covered call.

  5. none of the above.

At expiration, a profit is realized if the stock price is:

  1. between $52 and $68.

  2. below $60.

  3. above $60.

  4. below $52 or above $68.

  5. none of the above.

Before expiration, the time value of a call option is equal to:

  1. zero.

  2. the actual call price minus the intrinsic value of the call.

  3. the intrinsic value of the call.

  4. the actual call price plus the intrinsic value of the call.

  5. none of the above.
Answer #1

1.
a short straddle [A short straddle is writing 1 call and writing 1 put]

2.
Profit=-MAX(S-60,0)-MAX(60-S,0)+5+3
S>60: Profit=-S+68
S<60: Profit=S-52

Profit is realized between $52 and $68.

3.
the actual call price minus the intrinsic value of the call.